OpenStudy (anonymous):

what is the equation for the inverse of y=cos x+1?

OpenStudy (danjs):

\[y = \cos(x) + 1\] switch y--x around and solve for y, that is the inverse function

OpenStudy (danjs):

\[x = \cos(y) +1\]

OpenStudy (danjs):

solve for y=

OpenStudy (anonymous):

y=1? @DanJS

OpenStudy (anonymous):

so y=1-x ? @DanJS

OpenStudy (danjs):

The inverse function, you have to change variables x and y, then solve for y x = cos(y) + 1 solving for y \[y = \cos^{-1}(x-1)\]

OpenStudy (anonymous):

x=2,y=0 @DanJS